# UoR-NCL at SemEval-2025 Task 1: Using Generative LLMs and CLIP Models for Multilingual Multimodal Idiomaticity Representation

### v2

@@ Abstract
SemEval-2025 Task 1 focuses on ranking images based on their alignment with a given nominal compound that may carry idiomatic meaning in both English and Brazilian Portuguese. To address this challenge, this work uses generative large language models (LLMs) and multilingual CLIP models to enhance idiomatic compound representations. LLMs generate idiomatic meanings for potentially idiomatic compounds, enriching their semantic interpretation. These meanings are then encoded using multilingual CLIP models, serving as representations for image ranking. Contrastive learning and data augmentation techniques are applied to fine-tune these embeddings for improved performance. Experimental results show that multimodal representations extracted through this method outperformed those based solely on the original nominal compounds. The fine-tuning approach shows promising outcomes but is less effective than using embeddings without fine-tuning. The source code used in this paper is available at https://github.com/tongwu17/SemEval-2025-Task1-UoR-NCL.
@@
Thanet Markchom1and Tong Wu2and Liting Huang3and Huizhi Liang3 1Department of Computer Science, University of Reading, Reading, UK 2 Formerly at School of Computing, Newcastle University, Newcastle upon Tyne, UK 3School of Computing, Newcastle University, Newcastle upon Tyne, UK thanet.markchom@reading.ac.uk, tongwuwhitney@gmail.com , L.Huang29@newcastle.ac.uk, huizhi.liang@newcastle.ac.uk
@@ 1 Introduction
Existing NLP models, particularly those based on transformer architectures such as GPT Radford and Narasimhan (2018) and CLIP (Contrastive Language–Image Pre-training) Radford et al. (2021), have made significant strides in language representation Markchom et al. (2022); Phelps et al. (2024); Xiong et al. (2024). However, they often struggle with idiomatic expressions due to their reliance on surface-level word associations and compositional semantics He et al. (2024). This problem necessitates further exploration of methods that can improve the models’ capacity to understand and represent idioms effectively.
@@ 2.3 Ensemble Method
When generating idiomatic meanings for compounds, multiple LLMs can be utilized to capture diverse interpretations. To further enhance image ranking, an ensemble approach leveraging multiple LLMs is proposed. For each input (a compound, an image, and a caption), each LLM generates its interpretation of the compound’s idiomatic meaning. A ranking score for the images is then computed based on these meanings. The individual scores from the LLMs are averaged to produce a final ranking score for each image. This ensemble method integrates insights from multiple LLMs, thereby improving the overall ranking performance.
@@ 4.2 Image Ranking Results
Based on the results of CI and CIC, LABSE-14 demonstrated the highest effectiveness in ranking. Consequently, the embeddings obtained using LABSE-14 with different LLMs were fine-tuned in CI-F and CIC-F. Multiple contrastive models were trained on individual sets of embeddings from various LLMs. The selected hyperparameters for each model can be found in Appendix A. Overall, the fine-tuned embeddings did not perform as well as the non-fine-tuned embeddings. Figure 2 shows the training and validation losses, as well as the test accuracy, during the fine-tuning of embeddings obtained using LABSE-14 with GPT-3.5, GPT-4, and GPT-4o. These figures suggest that the models effectively learned the fine-tuned embeddings, as test accuracy gradually increased over training epochs. However, the models began overfitting before the test accuracy could improve further. This could be due to the amount of training data being insufficient for the model to generalize well to unseen data. Due to the lack of performance improvement on the English dataset during fine-tuning, experiments on the Portuguese dataset were not conducted.
@@
Figure 2: Training loss, validation loss, and test accuracy, during the fine-tuning of embeddings obtained using LABSE-14 with GPT-3.5, GPT-4, and GPT-4o.
@@ 5 Conclusions
This work explored the use of generative LLMs and multilingual CLIP models to enhance idiomatic compound representations for image ranking in SemEval-2025 Task 1. By using LLMs to generate idiomatic meanings and leveraging multilingual CLIP models to extract multimodal embeddings, the proposed method improved representation quality compared to using original nominal compounds. Experimental results demonstrated the effectiveness of the proposed method. For English, the ensemble method using GPT-3.5, GPT-4, and GPT-4o, with the XLM-R Large ViT-B/32 multilingual CLIP model achieved superior performance compared to the other selected LLMs and CLIP models. For Brazilian Portuguese, GPT-3.5 with the LABSE ViT-L/14 multilingual CLIP model outperformed the others. Fine-tuning CLIP embeddings performed worse than using embeddings extracted from pretrained CLIP models. This is likely due to limitations in fine-tuning data and the capacity of the proposed contrastive learning model. However, it could still be a promising approach for further improvement. Future work could focus on refining fine-tuning strategies and expanding training data to further enhance idiomaticity representation.

### v3

@@ Abstract
SemEval-2025 Task 1 focuses on ranking images based on their alignment with a given nominal compound that may carry idiomatic meaning in both English and Brazilian Portuguese. To address this challenge, this work uses generative large language models (LLMs) and multilingual CLIP models to enhance idiomatic compound representations. LLMs generate idiomatic meanings for potentially idiomatic compounds, enriching their semantic interpretation. These meanings are then encoded using multilingual CLIP models, serving as representations for image ranking. Contrastive learning and data augmentation techniques are applied to fine-tune these embeddings for improved performance. Experimental results show that multimodal representations extracted through this method outperformed those based solely on the original nominal compounds. The fine-tuning approach shows promising outcomes but is less effective than using embeddings without fine-tuning.
@@
Thanet Markchom1and Tong Wu2and Liting Huang3and Huizhi Liang3 1Department of Computer Science, University of Reading, Reading, UK 2 Previously at School of Computing, Newcastle University, Newcastle upon Tyne, UK 3School of Computing, Newcastle University, Newcastle upon Tyne, UK thanet.markchom@reading.ac.uk, tongwuwhitney@gmail.com , L.Huang29@newcastle.ac.uk, huizhi.liang@newcastle.ac.uk
@@ 1 Introduction
Existing NLP models, particularly those based on transformer architectures such as GPT Radford et al. (2018) and CLIP (Contrastive Language–Image Pre-training) Radford et al. (2021), have made significant strides in language representation Markchom et al. (2022); Phelps et al. (2024); Xiong et al. (2024). However, they often struggle with idiomatic expressions due to their reliance on surface-level word associations and compositional semantics He et al. (2024). This problem necessitates further exploration of methods that can improve the models’ capacity to understand and represent idioms effectively.
@@ 2.3 Ensemble Method
When generating idiomatic meanings for compounds, multiple LLMs can be utilized to capture diverse interpretations. To further enhance image ranking, an ensemble approach leveraging multiple LLMs is proposed. For each input (a compound, an image, and a caption), each LLM generates its interpretation of the compound’s idiomatic meaning. A ranking score for the images is then computed based on these meanings. The individual scores from the LLMs are averaged to produce a final ranking score for each image. There is no weighting, i.e., each LLM contributes equally to the final ranking score. As for consistency, each model may interpret idiomatic meanings slightly differently and may not always be consistent with others. However, it is assumed that the majority of models will converge on the correct interpretation. By averaging their scores, individual biases are smoothed out, and commonly accurate interpretations are reinforced. Overall, this ensemble method integrates insights from multiple LLMs, thereby improving the overall ranking performance.
@@ 4.2 Image Ranking Results
Based on the results of CI and CIC, LABSE-14 demonstrated the highest effectiveness in ranking. Consequently, the embeddings obtained using LABSE-14 with different LLMs were fine-tuned in CI-F and CIC-F. Multiple contrastive models were trained on individual sets of embeddings from various LLMs. The selected hyperparameters for each model can be found in Appendix A. Overall, the fine-tuned embeddings did not perform as well as the non-fine-tuned embeddings. Figure 2 shows the training and validation losses, as well as the test accuracy, during the fine-tuning of embeddings obtained using LABSE-14 with GPT-3.5, GPT-4, and GPT-4o. These figures suggest that the models effectively learned the fine-tuned embeddings, as test accuracy gradually increased over training epochs. However, the models began overfitting before the test accuracy could improve further. This could be due to the amount of training data being insufficient for the model to generalize well to unseen data. Extra data augmentation could improve fine-tuning by introducing linguistic and visual diversity within existing, seen idioms. This may help the model more accurately match images in different styles to the generated meanings of seen idioms expressed with varying wordings. However, this may not be effective for unseen idioms if they share no common meanings with those in the training set. Similarly, the use of regularization may help generalize the model’s ability to match images with seen idioms in the training set, but it might not improve generalization to unseen idioms. Due to the lack of performance improvement on the English dataset during fine-tuning, experiments on the Portuguese dataset were not conducted.
@@
Figure 2: Training loss, validation loss, and test accuracy, during the fine-tuning of embeddings obtained using LABSE-14, with GPT-3.5, GPT-4, and GPT-4o used for idiomatic meaning generation.
@@
(a) $K$
(b) $\tau$
(c) Dropout rate
Figure 3: Top-1 accuracy on the test set during the fine-tuning of embeddings obtained using LABSE-14, with GPT-3.5 used for idiomatic meaning generation.
(d) $K$
(e) $\tau$
(f) Dropout rate
Figure 4: Top-1 accuracy on the test set during the fine-tuning of embeddings obtained using LABSE-14, with GPT-4 used for idiomatic meaning generation.
(a) $K$
(b) $\tau$
(c) Dropout rate
Figure 5: Top-1 accuracy on the test set during the fine-tuning of embeddings obtained using LABSE-14, with GPT-4o used for idiomatic meaning generation.
4.3 Hyperparameter Analysis
This section explores the impact of key hyperparameters on model fine-tuning performance, including the number of soft negatives ($K$), temperature ($\tau$), and dropout rate. For each fine-tuned model, one hyperparameter was varied while the others remained at their optimal values (as in Table 3). Top-1 accuracy was evaluated across training epochs to assess each hyperparameter. Figures 5, 5, and 5 illustrate the effect and sensitivity of three hyperparameters on the top-1 accuracy during the fine-tuning of LABSE-14 embeddings, using different GPT variants for idiomatic meaning generation. Across all models, $K$ exhibited moderate sensitivity, with higher $K$ generally yielding better results for GPT-3.5 and GPT-4. Meanwhile, for GPT-4o, lower $K$ generally performed better. Temperature $\tau$ showed high sensitivity, with small variations (from 0.08 to 0.1) leading to notable shifts in accuracy. For GPT-3.5 and GPT-4, $\tau=0.1$ yielded the best results, whereas for GPT-4o, a lower value of $\tau=0.08$ was optimal. The dropout rate exhibited model-specific effects. A rate of $0.1$ worked best for GPT-3.5. For GPT-4, lower rates of $0.1$ and $0.3$ yielded similar performance. For both GPT-3.5 and GPT-4, higher dropout rates appeared to degrade performance. In contrast, GPT-4o benefited from a higher rate of $0.5$.
@@ 5 Conclusions
This work explored the use of generative LLMs and multilingual CLIP models to enhance idiomatic compound representations for image ranking in SemEval-2025 Task 1. By using LLMs to generate idiomatic meanings and leveraging multilingual CLIP models to extract multimodal embeddings, the proposed method improved representation quality compared to using original nominal compounds. Experimental results demonstrated the effectiveness of the proposed method. For English, the ensemble method using GPT-3.5, GPT-4, and GPT-4o, with the XLM-R Large ViT-B/32 multilingual CLIP model achieved superior performance compared to the other selected LLMs and CLIP models. For Brazilian Portuguese, GPT-3.5 with the LABSE ViT-L/14 multilingual CLIP model outperformed the others. Fine-tuning CLIP embeddings performed worse than using embeddings extracted from pretrained CLIP models. This is likely due to limitations in fine-tuning data and the capacity of the proposed contrastive learning model. However, it could still be a promising approach for further improvement. Future work could focus on improving caption utilization (e.g., through different truncation methods and paraphrasing), refining fine-tuning strategies and expanding training data to further enhance idiomaticity representation.